Apparent Weight You take a ride in a fast
elevator to the top of a tall building and ride
back down. Compare your apparent and real
weights at each part of the journey. Sketch
free-body diagrams to support your answers.
1
Expert's answer
2020-03-31T09:32:21-0400

Consider two cases , where the fast elevator is going up and going down.




For each of the cases, apparent weight is given by R as shown in the free body diagram.


For the going upward case,

From the Free body diagram.


"F=ma\\
R_1 -mg= ma \\
R_1=m(g+a)"


Therefore the apparent weight will be greater than the real weight if the elevator is going upwards and accelerating "(a>0)" and will be lower than the real weight if the elevator is decelerating (a<0).


For the going downward case,

From the Free body diagram.


"F=ma\\
mg-R_2=ma\\
R_2=m(g-a)\\"


Therefore the apparent weight will be lower than the real weight if the elevator is going downwards and accelerating "(a>0)" and will be higher than the real weight if the elevator is decelerating "(a<0)" .